OpenStudy (anonymous):

Two jets are beginning their flight at the Louisville International Airport. One starts at 1:00 p.m. and flies east at 395 mph. Another starts one hour later and flies west at 400 mph. In how many hours after the first plane left, will they be 2,000 miles apart? Round to the nearest tenth of an hour. A. 2.0 B. 3.0 C. 2.5 D. 3.5 Please Explain!

OpenStudy (anonymous):

RT+R(T-1)=D 395x+400(x-1)=2000 395x+400x-400=2000 795x=2400 x=3.018 hours Answer-->Rounded off, it will be 3.0 hours after the first plane takes off.